Sunland Company purchased 100 of the 1000 outstanding shares of Darby Company's common stock for $660000 on January 2, 2018. During 2018, Darby Company declared dividends of $130000 and reported earnings for the year of $460000.

If Sunland Company uses the equity method of accounting for its investment in Darby Company, its Equity Investments (Darby) account at December 31, 2018 should be

$660000.
$693000.
$706000.
$647000.
Answer #1

Ownership percentage = 100/1000 = 10%

Equity investment account balance

= 660,000 + Share in earnings - Share in dividends

= 660,000 + (460,000*10%) - (130,000*10%)

= 660,000 + 46,000 - 13,000

= 693,000

Option B is the answer